OpenStudy (anonymous):

Sam is driving at a constant speed of 120km/h and is following Alex's vehicle at a safe distance (3s). Just as Alex passes a telephone pole, the pole topples across the road. Same slams down on the brakes and comes to a stop being smashing into the pole. How far away from the telephone pole was Same when it fell across the road? What was Sam's acceleration How long did it take for him to stop?

OpenStudy (anonymous):

Show steps, not just the answer, thanks.

OpenStudy (anonymous):

what is the 3s?

OpenStudy (anonymous):

3 seconds

OpenStudy (anonymous):

what i am asking you is what does it represent?

OpenStudy (anonymous):

if you have the velocity he is travelling at and the time, you can solve for the distance. and also the acceleration

OpenStudy (anonymous):

v=s/t so t*v=s and a=v/t

OpenStudy (anonymous):

it represents that it's 3 seconds away from the other vehicle. i'm not understanding your equations.